# Large Language Model-assisted Autonomous Vehicle Recovery from Immobilization

### v1

@@ IV Experiment
In this section, we employ the CARLA simulator [17] (version 0.9.15.1) for closed-loop driving performance evaluation, including both quantitative and qualitative analyses. The quantitative evaluation is conducted using the Bench2Drive benchmark  [35], while the qualitative evaluation is performed on scenarios from the training dataset provided by Bench2Drive.
TABLE I: Results on Bench2Drive Closed-loop Benchmark
Method Venue Scheme Modality Closed-loop Metrics
DS $\uparrow$ SR(%) $\uparrow$ Efficiency $\uparrow$ Comfort $\uparrow$
TCP-traj NeurIPS 2022 IL Image 59.90 30.00 76.54 18.08
AD-MLP arXiv 2023 IL State 18.05 0.00 48.45 22.63
UniAD-Base CVPR 2023 IL Image 45.81 16.36 129.21 43.58
VAD ICCV 2023 IL Image 42.35 15.00 157.94 46.01
ThinkTwice CVPR 2023 IL Image 62.44 31.23 69.33 16.22
DriveAdapter ICCV 2023 IL Image 64.23 30.22 70.22 16.01
GenAD ECCV 2024 IL Image 44.81 15.90
DriveTrans ICLR 2025 IL Image 63.46 35.01 100.64 20.78
MomAD CVPR 2025 IL Image 44.54 16.71 170.21 48.63
Raw2Drive (SOTA) arXiv 2025 RL Image 71.36 50.24 214.17 22.42
Behavior agent Image, Lidar 48.70 18.21 68.37 43.68
Behavior agent + StuckSolver (Ours) Zero-shot Image, Lidar 65.23 36.32 87.49 46.74
Behavior agent + StuckSolver (Ours) Zero-shot Image, Lidar 70.89 50.01 87.78 47.31
IL: imitation learning; RL: reinforcement learning. denotes expert feature distillation. indicates the use of human guidance in certain scenarios.
The Behavior agent, provided by CARLA, does not include a perception module.
We extend it by integrating a perception module equipped with sensors such as cameras and LiDAR.
IV-A Quantitative Evaluation
@@
Benchmark and Metrics. We use Bench2Drive  [35], a closed-loop evaluation protocol under CARLA Leaderboard 2.0 for end-to-end autonomous driving. It includes 220 short routes with one challenging corner case per route for analysis of different driving abilities. We employ the official metrics of Bench2Drivev for evaluation. Infraction Score (IS) measures the number of infractions made along the route, accounting for interactions with pedestrians,vehicles, road layouts, red lights,etc. Route Completion (RC) indicates the percentage of the route completed by the autonomous agent. Driving Score (DS),is calculated as the product of Route Completion and Infraction Score. Success Rate (SR) is used, which measures the proportion of successfully completed routes within the allotted time and without traffic violations. Efficiency checks whether the AV’s speed is too low. This is determined by comparing the AV’s speed with nearby vehicle. Comfortness measures how closely the AV’s motion smoothness matches that of human drivers by checking whether dynamic variables like acceleration, yaw rate, and jerk stay within expert-derived comfort bounds.
Results. From the results in Table 1, the rule-based behavior agent shows limited performance due to its relatively simple decision-making module, which often becomes stuck in challenging corner cases within the Bench2Drive benchmark, leading to low DS and SR scores. Moreover, because the behavior agent remains stationary after getting stuck, creating a large speed gap between itself and surrounding traffic agents, its efficiency metric is also poor. When integrated with StuckSolver, however, the system achieves consistent improvements across all metrics, with particularly notable gains in DS and SR. Its performance ranks second only to the SOTA end-to-end method. The lower efficiency compared to SOTA is mainly because StuckSolver initiates reasoning and recovery only after immobilization is detected, causing the vehicle to pause for several seconds before resuming motion once a recovery plan is generated. For the efficiency and comfort metrics, although both show improvement, the gains are less pronounced compared to DS and SR. This is because StuckSolver only intervenes after the vehicle has remained at low speed or stationary for a certain duration. As a result, the AV’s speed often falls below the average speed of surrounding vehicles but rarely exceeds it, limiting efficiency gains. In addition, the reasoning time required by StuckSolver also contributes to this effect.
Moreover, when human guidance is provided in certain scenarios, the system achieves higher performance (DS: 70.89, SR: 50.01), coming very close to the SOTA results. These results highlight the strength of our approach: without any fine-tuning or architectural modification, simply integrating StuckSolver with a relatively simple rule-based AV system yields performance comparable to that of sophisticated end-to-end methods trained with extensive data and fine-tuning. Furthermore, the system’s performance can be further enhanced when human guidance is incorporated.
Figure 4: Stucksolver recover the immobilized AV in construction and parked obstacle scenarios. The green line is the re-planned trajectory. The green and blue bounding boxes indicate the regions that the AV and surrounding vehicles are expected to traverse within the next few seconds.
Figure 5: Experiment results of the vehicle opens door (two ways) scenario and the pedestrian crossing scenario. The upper part shows the RGB images captured by the ego vehicle’s front-view camera, while the lower part illustrates the decision-making process of the StuckSolver.
In addition, we further analyze the cases where the AV fails before integrating StuckSolver but successes afterward. Since the AV is controlled by a rule-based decision module combined with low-latency control algorithms, it performs well in time-sensitive scenarios (e.g., hard braking, pedestrian crossings) as well as in regular traffic situations (e.g., signalized junction left turn). However, because the decision module only drives the vehicle along the pre-planned route without triggering replanning, the AV fails in scenarios such as construction and parked obstacle, where it simply stops, leading to reduced DS, SR, and efficiency. With StuckSolver integrated, once the AV stops for longer than $\delta t=1s$, the vehicle is identified as immobilized. At this point, StuckSolver initiates scene analysis and generates a recovery plan to restore normal operation, as shown in Fig. 4.
IV-B Qualitative Evaluation
To evaluate the rationality of StuckSolver’s decision-making process, two representative scenarios from the Bench2Drive training dataset were selected: the vehicle opens door (two-way) scenario and the pedestrian crossing scenario.
In the vehicle opens door (two-way) scenario, a stationary white sedan is located in the adjacent right lane with its left door fully open. After observation, StuckSolver correctly determines that the ego vehicle is stuck because the open door of the white vehicle blocks its current lane. It then reasonably decides to perform a left lane change to avoid the obstruction. In the pedestrian crossing scenario, StuckSolver identifies that the ego vehicle has stopped due to a red traffic light and pedestrian crossing. Consequently, it determines that the vehicle is not stuck and appropriately chooses not to intervene, outputting no behavior plan.
@@ V Conclusion and Future Work
In this paper, we propose StuckSolver, an LLM-based recovery framework designed to enhance AV autonomy by enabling self-recovery from immobilized situations. Unlike traditional recovery methods that rely on remote intervention or manual takeover, StuckSolver leverages LLMs’s advantages to detect and resolve immobilization scenarios or interpret passenger guidance when autonomous recovery is insufficient. Designed as a seamlessly integrable add-on module, StuckSolver requires minimal modifications to existing AV systems. We evaluate StuckSolver on the Bench2Drive benchmark as well as in the scenarios of the Bench2Drive dataset. Results on the benchmark show that, without any passenger guidance, StuckSolver achieves performance close to the SOTA. When passenger guidance is incorporated in certain scenarios, StuckSolver’s performance is further enhanced. Additionally, in the scenarios of the Bench2Drive dataset, StuckSolver can accurately determine whether the AV is truly stuck. When a stuck situation occurs, it is capable of taking reasonable actions to help the AV escape, thereby improving the reliability, and operational efficiency of AVs.
Looking ahead, future work will focus on improving the inference efficiency of StuckSolver. At present, its average inference time is 2.8 seconds per query, which prevents it from handling certain time-sensitive scenarios. To address this, we plan to distill a lightweight yet capable LLM that offers significantly faster inference speed. In addition, our current approach is designed for rule-based or modular AV systems and is not directly applicable to end-to-end AV frameworks. In the future, we intend to explore integration strategies that enable StuckSolver to complement and enhance end-to-end autonomous driving systems.

### v2

@@ IV Experiment
In this section, we conduct closed-loop driving performance evaluations using the Bench2Drive benchmark [35] within the CARLA simulator [17] (version 0.9.15). The experiments aim to quantitatively assess the effectiveness of StuckSolver in recovering immobilized autonomous vehicles. StuckSolver is built on GPT-4o-2024-08-06. All evaluations are performed on a workstation running Ubuntu 22.04, equipped with an AMD Ryzen Threadripper 7960X (24 cores) CPU and two NVIDIA RTX 5000 Ada GPUs.
@@
Benchmark and Metrics. The Bench2Drive benchmark [35] under CARLA Leaderboard 2.0 consists of 220 short routes, each containing one challenging corner case designed to evaluate different driving abilities. We adopt the official Bench2Drive metrics for evaluation. Infraction Score (IS) measures the number of violations committed along a route, including interactions with pedestrians, vehicles, road layouts, and red lights. Route Completion (RC) represents the percentage of the route completed by the autonomous agent. Driving Score (DS) is calculated as the product of Route Completion and Infraction Score. Success Rate (SR) measures the proportion of routes completed within the allotted time without traffic violations. Efficiency evaluates whether the AV’s speed is excessively low by comparing it with surrounding vehicles. Comfort assesses motion smoothness by checking whether dynamic variables such as acceleration, yaw rate, and jerk remain within expert-defined comfort thresholds that represent human driving behavior.
Results. The simulation results in Table 1 show that our method, which integrates StuckSolver with the CARLA rule-based Behavior Agent, achieves near state-of-the-art performance in DS (65.23), SR (36.32), and Comfort (46.74). When combined with passenger guidance, the performance further improves to DS 70.89, SR 50.01, and Comfort 47.31, demonstrating the adaptability and robustness of the proposed framework. During simulation, we continuously monitor the AV’s behavior and provide passenger instructions when we consider them beneficial, such as suggesting a lane change to navigate around an obstacle. These human-in-the-loop interactions occur in about 15% of the 220 routes. The provided passenger guidance is not necessarily optimal and represents just one possible form of human input. Our intent is not to optimize human intervention but to demonstrate that incorporating human guidance, even in a simple and limited form, can effectively improve AV decision-making and overall driving performance.
In terms of efficiency, the Behavior Agent with StuckSolver performs below some learning based baselines(e.g., VAD, UniAD-Base, and MomAD). This difference primarily arises from the simplicity of the CARLA rule-based agent, which lacks continuous optimization and relies on discrete heuristic decision rules. Moreover, StuckSolver is activated only when immobilization risks are detected, unlike end-to-end foundation models that operate continuously and incur high computational costs. As a result, the rule-based agent may occasionally execute suboptimal maneuvers, such as missing opportunities to change to a faster lane to catch up in speed.
Overall, the results highlight a fundamental tradeoff between model complexity and operational efficiency. End-to-end approaches achieve strong performance but require substantial computation, whereas lightweight rule-based methods are efficient but less adaptive. StuckSolver bridges these two ends of the spectrum by enhancing simple and interpretable driving agents without significantly increasing computational burden. By adjusting its activation frequency through parameters such as the minimum speed threshold $v_{min}$ and the time interval $\triangle t$, the framework can shift between these extremes. A higher activation frequency makes the system behave more like an end-to-end model with the LLM engaged almost continuously, while a lower frequency allows the rule-based agent to dominate. This tunable design provides a practical means to balance computational effort and solution quality for efficient and adaptive autonomous driving.
To evaluate the rationality of StuckSolver’s decision-making process, three representative scenarios from the Bench2Drive training dataset were analyzed. In the pedestrian crossing scenario, StuckSolver correctly identifies that the ego vehicle has stopped for a red traffic light and pedestrians. It determines that the vehicle is not immobilized and appropriately refrains from intervention, producing no behavior plan. In the unexpected door-opening scenario, a stationary white sedan in the adjacent right lane has its left door fully open. After observation, StuckSolver recognizes that the ego vehicle is indeed stuck because the open door blocks its current lane. It then reasonably decides to execute a left lane change to avoid the obstruction. In the cyclist following scenario, the autonomous vehicle is moving slowly behind two cyclists. Upon receiving a verbal instruction from the passenger to overtake, StuckSolver evaluates the surrounding traffic conditions, deems the maneuver safe, and generates a left lane change behavior plan that enables the vehicle to pass the cyclists smoothly and safely.
@@ V Conclusion and Future Work
In this paper, we present StuckSolver, an LLM-based recovery framework designed to enhance AV autonomy by enabling self-recovery from immobilized situations. Unlike traditional recovery methods that depend on remote intervention or manual takeover, StuckSolver leverages the reasoning capability of LLMs to detect and resolve immobilization scenarios and interpret passenger guidance when autonomous recovery alone is insufficient. Designed as a seamlessly integrable add-on module, StuckSolver connects to existing AV systems through structured APIs, reading and writing data without modifying any internal modules or control logic. We evaluate its effectiveness on the Bench2Drive benchmark by integrating it with the rule-based Behavior Agent in CARLA. Results show that, even without passenger guidance, StuckSolver achieves DS and SR performance close to state-of-the-art end-to-end models, and that its performance further improves when passenger guidance is incorporated in specific scenarios.
Looking ahead, future work will focus on enhancing the inference efficiency of StuckSolver. Currently, its average inference time is about 2.8 seconds per query, which limits its ability to respond to time-sensitive scenarios. To address this, we plan to develop a distilled version of the underlying LLM that retains strong reasoning capability while significantly reducing latency. Another important direction is real-world testing and deployment under edge computing environments to evaluate StuckSolver’s robustness, responsiveness, and scalability in practical AV operations. In addition, we plan to further explore adaptive interaction strategies that enable StuckSolver to better interpret and learn from human guidance over time, improving its contextual understanding and decision-making in complex driving environments.